# Self-Distilled Reasoner: On-Policy Self-Distillation for Large Language Models

### v1

@@ Abstract
Knowledge distillation improves large language model (LLM) reasoning by compressing the knowledge of a teacher LLM to train smaller LLMs. On-policy distillation advances this approach by having the student sample its own trajectories while a teacher LLM provides dense token-level supervision, addressing the distribution mismatch between training and inference in off-policy distillation methods. However, on-policy distillation typically requires a separate, often larger, teacher LLM and does not explicitly leverage ground-truth solutions available in reasoning datasets. Inspired by the intuition that a sufficiently capable LLM can rationalize external privileged reasoning traces and teach its weaker self (i.e., the version without access to privileged information), we introduce On-Policy Self-Distillation (OPSD), a framework where a single model acts as both teacher and student by conditioning on different contexts. The teacher policy conditions on privileged information (e.g., verified reasoning traces) while the student policy sees only the question; training minimizes the per-token divergence between these distributions over the student’s own rollouts. We demonstrate the efficacy of our method on multiple mathematical reasoning benchmarks, achieving 4-8× token efficiency compared to reinforcement learning methods such as GRPO and superior performance over off-policy distillation methods.
@@ 1 Introduction
Figure 1: Overview of On-Policy Self-Distillation (OPSD): Given a reasoning dataset $\mathcal{S}=\{(x_{i},y_{i}^{\star})\}_{i=1}^{N}$, we instantiate two policies from the same LLM: a student policy $p_{S}(\cdot\mid x)$ and a teacher policy $p_{T}(\cdot\mid x,y^{\star})$. The student generates an on-policy response $\hat{y}\sim p_{S}(\cdot\mid x)$. Both policies then evaluate this trajectory to produce next-token distributions $p_{S}(\cdot\mid x,\hat{y}_{<n})$ and $p_{T}(\cdot\mid x,y^{\star},\hat{y}_{<n})$ at each step $n$. The learning objective minimizes the per-token divergence $D(p_{T}\|p_{S})$ along the student’s rollout. Crucially, gradients backpropagate only through the student’s logits, allowing the model to self-distil.
@@
  • We evaluate OPSD on four competition-level mathematical reasoning tasks, demonstrating that it outperforms both RLVR (e.g., GRPO) and supervised fine-tuning baselines.
  • We show that OPSD achieves better performance with nearly $8\times$ improved token efficiency and lower computational cost than GRPO.
@@ 3 Methods
Figure 2: Prompt example for student and teacher policies. Both policies share the same parameters $\theta$ but differ in conditioning context. The teacher receives the ground-truth solution $y^{\star}$ as privileged information before generation. To ensure a natural transition before evaluating the student’s rollout, the teacher is prompted to rationalize and generate its own solution.
@@ 3.1 Learning from Verifiable Reasoning Dataset
Reasoning dataset $\mathcal{S}=\{(x_{i},y_{i}^{\star})\}_{i=1}^{N}$; language model $p_{\theta}$; divergence DD(e.g., $\mathrm{JSD}_{\beta}$) Define student policy $p_{S}(\cdot\mid x):=p_{\theta}(\cdot\mid x)$Define teacher policy $p_{T}(\cdot\mid x,y^{\star}):=p_{\theta}(\cdot\mid x,y^{\star})$same parameters; different conditioning not converged Sample a minibatch $\mathcal{B}\subset\mathcal{S}$$(x,y^{\star})\in\mathcal{B}$Sample on-policy response $\hat{y}\sim p_{S}(\cdot\mid x)$Compute the token-wise divergence along the student rollout:
@@
[…]
@@ 3.2 On-Policy Self-Distillation
Critically, both policies share the same parameters $\theta$ but differ only in their conditioning context. The teacher has access to information unavailable at test time, allowing it to provide informed guidance. To encourage the teacher to naturally evaluate the student’s generation, we add a prompt asking the teacher to generate a new solution after rationalization, as shown in Figure 2.
@@
Alternatively, following recent on-policy distillation methods (lu2025onpolicydistillation), we form a sampled-token shaping signal (equivalently, a reverse-KL signal on sampled actions) and optimize with policy gradient. For each position $n$ in a sampled sequence $\hat{y}$, define the advantage term
@@
Table 2: Performance comparison across mathematical reasoning benchmarks for Qwen3 models from 1.7B to 8B. We report average@16 using suggested sampling parameters from the Qwen3 blog with temperature of 1.2 and generation length of 38k, with detailed parameter in Table 5.
@@ 4 Experiments
In this section, we conduct comprehensive experiments to answer the following research questions:
@@
    How does OPSD compare to SFT and GRPO in terms of mathematical reasoning performance and what’s the improved sample efficiency? (§4.2)
@@
    How does OPSD scale across different model sizes, does self-distillation require more powerful model ability? (§4.3.1)
@@
    What is the effect of generation length on training performance and sample efficiency? (§4.3.2)
@@
    Does computing divergence over the full vocabulary logits provide benefits compared to computing it only over sampled tokens and optimizing through policy gradient? (§4.3.3)
@@ 4.1 Experimental Setup
Baselines. We compare against two methods trained on the same dataset: (1) SFT, standard supervised fine-tuning on expert trajectories, which can be seen as off-policy distillation from a more powerful LLM that generated the reasoning traces; (2) GRPO (shao2024deepseekmath), group relative policy optimization with binary outcome rewards verified against ground-truth answers.
Implementation details. For GRPO, we sample 8 responses per problem. For OPSD, we sample 1 response per problem. We use Adam optimizer with a learning rate of 1e-5, warmup ratio of 0.1, and cosine learning rate decay. For the divergence measure in Eq. 6, we use $\operatorname{JSD}_{\beta=0.5}$. Importantly, we fix the teacher policy to be the initial policy, rather than the currently updating learning policy, as we find this helps stabilize training and implicitly acts as regularization to prevent excessive deviation from the initial policy. All experiments are conducted on 8×A100 GPUs with LoRA (hu2022lora). More experimental details are in Appendix 8.1.
@@ 4.2 Main Results
[…]
Table 2 reports results on competition-level mathematical reasoning benchmarks. OPSD consistently outperforms SFT and improves over the base model across scales; it matches or exceeds GRPO at 4B/8B, and is comparable at 1.7B. Notably, OPSD accomplishes these gains using only a single rollout per problem, whereas GRPO requires 8 rollouts, demonstrating improved sample efficiency.
@@
In addition to improved accuracy, OPSD is significantly more token-efficient than GRPO. Figure 3 compares the two methods under the same effective training batch size on Qwen3-4B. While GRPO relies on 8 rollouts with long generation budgets of 16k, OPSD achieves higher performance using substantially fewer generated tokens of 2k and needs only 1 rollout per prompt. This efficiency stems from dense token-level supervision from the teacher distribution, reducing sampling cost and training time without sacrificing performance. We hypothesize that the early tokens are more important for distillation than the later tokens, as the earlier tokens can represent more important branching points.
@@ 4.3 Discussions
4.3.1 Effect of Model Scale
Our method relies on the teacher policy’s ability to rationalize reference solutions when conditioned on privileged information. Under a fixed dataset, this capability depends on sufficient model capacity and is expected to scale with model size. We therefore hypothesize that OPSD becomes increasingly effective as models grow more capable of leveraging privileged context. To evaluate this, we apply OPSD to the Qwen3 family at three scales: 1.7B, 4B, and 8B parameters. As shown in Table 2, OPSD provides limited gains over GRPO at the 1.7B scale although OPSD still improves over base and SFT at 1.7B., while yielding progressively larger improvements at the 4B and 8B scales, consistent with our hypothesis.
4.3.2 Effect of Generation Length
@@ 5 Related Work
Improving LLM Reasoning through SFT and RL. SFT and RL are two primary methods for improving LLM reasoning ability. SFT on high-quality reasoning traces has demonstrated strong performance (yu2023metamath; numina_math_datasets; pasteropenwebmath; openthoughts), and that smaller, carefully curated datasets can outperform larger but noisier collections (ye2025limoreasoning; muennighoff2025s1; zhou2023lima). However, prior work shows that SFT-based reasoning often relies on memorization rather than robust generalization (chu2025sft). In contrast, RL-based approaches optimize directly for outcome-based objectives can exhibit stronger transfer to novel problems (huan2025does). More recent algorithms such as GRPO (guo2025deepseek; shao2024deepseekmath) enable scalable RL by estimating advantages from group-level rewards without requiring an explicit critic as in PPO (schulman2017proximal). Building on this line of work, a growing body of research highlights the effectiveness of reinforcement learning with verifiable rewards (RLVR) for reasoning tasks (yu2025dapo; liu2025understanding; yue2025vapo; Polaris2025; zheng2025group).
LLM Self-Training.
Our work is related to a growing body of research demonstrating that LLMs can improve by generating and exploiting their own supervision signals (allentowards; xu2024survey; chen2024self). Self-Instruct (wang2023self) and Self-Align (sun2023principle) demonstrate that large language models can bootstrap instruction-following and alignment with minimal human supervision by leveraging small sets of human-written seeds—either instructions or principles—to generate synthetic training data. Context distillation (snell2022learning) shows that models can internalize the benefits of privileged context tokens (e.g., instructions or scratchpads) by training a student to reproduce the same outputs without access to such context at inference time through SFT. Recent work on in-context editing (qicontext) demonstrates that models can learn new knowledge by optimizing toward self-induced contextual distributions rather than one-hot targets for knowledge editing. In the reasoning domain, ReST (gulcehre2023reinforced) and STaR (zelikman2022star) improve performance through iterative loops of rationale generation, filtering based on rewards or ground-truth answers, and fine-tuning on successful samples. LLM can also be used as a judge to generate RL rewards (yuan2024self) for itself. While aligned with this self-training paradigm, OPSD introduces a distinct approach: we perform on-policy, token-level self-distillation where the model learns from its own outputs conditioned on privileged access to ground-truth solutions. This transforms reasoning improvement into learning a conditional distribution induced by both the dataset’s ground-truth answers and the model’s own understanding of how to reach them.
@@ 6 Conclusion
7 Limitations and Future Directions

### v2

@@ Abstract
Knowledge distillation improves large language model (LLM) reasoning by compressing the knowledge of a teacher LLM to train smaller LLMs. On-policy distillation advances this approach by having the student sample its own trajectories while a teacher LLM provides dense token-level supervision, addressing the distribution mismatch between training and inference in off-policy distillation methods. However, on-policy distillation typically requires a separate, often larger, teacher LLM and does not explicitly leverage ground-truth solutions available in reasoning datasets. Inspired by the intuition that a sufficiently capable LLM can rationalize external privileged reasoning traces and teach its weaker self (i.e., the version without access to privileged information), we introduce On-Policy Self-Distillation (OPSD), a framework where a single model acts as both teacher and student by conditioning on different contexts. The teacher policy conditions on privileged information (e.g., verified reasoning traces) while the student policy sees only the question; training minimizes the per-token divergence between these distributions over the student’s own rollouts. We demonstrate the efficacy of our method on multiple mathematical reasoning benchmarks, achieving 8-12× token efficiency compared to reinforcement learning methods such as GRPO and superior performance over off-policy distillation methods. Code repo: https://github.com/siyan-zhao/OPSD.
@@ 1 Introduction
Figure 1: Overview of On-Policy Self-Distillation (OPSD): Given a reasoning dataset $\mathcal{S}=\{(x_{i},y_{i}^{\star})\}_{i=1}^{N}$, we instantiate two policies from the same LLM: a student policy $p_{S}(\cdot\mid x)$ and a teacher policy $p_{T}(\cdot\mid x,y^{\star})$. The student generates an on-policy response $\hat{y}\sim p_{S}(\cdot\mid x)$. Both policies then evaluate this trajectory to produce next-token distributions $p_{S}(\cdot\mid x,\hat{y}_{<n})$ and $p_{T}(\cdot\mid x,y^{\star},\hat{y}_{<n})$ at each step $n$. The learning objective minimizes the per-token divergence $D(p_{T}\|p_{S})$ along the student’s rollout. The divergence here can be forward KL, reverse KL or JSD. Crucially, gradients backpropagate only through the student’s logits, allowing the model to self-distil.
@@
  • We evaluate OPSD on four competition-level mathematical reasoning tasks, demonstrating that it matches the performance of GRPO and outperform supervised fine-tuning baselines.
  • We show that OPSD achieves up to $12\times$ improved token efficiency than GRPO.
@@ 3 Methods
Figure 2: Prompt example for student and teacher policies. Both policies share the same parameters $\theta$ but differ in conditioning context. The teacher receives the ground-truth solution $y^{\star}$ as privileged information before generation. To ensure a natural transition before evaluating the student’s rollout, the teacher is prompted to rationalize and generate its own solution. Note that the teacher won’t be generating tokens—rationalization is done implictly through one forward pass.
@@ 3.1 Learning from Verifiable Reasoning Dataset
Reasoning dataset $\mathcal{S}=\{(x_{i},y_{i}^{\star})\}_{i=1}^{N}$; language model $p_{\theta}$; divergence DD(e.g., $\mathrm{JSD}_{\beta}$) Let $p_{S}(\cdot\mid x)$and $p_{T}(\cdot\mid x,y^{\star})$be the same model $p_{\theta}$under different conditioning. not converged Sample a minibatch $\mathcal{B}\subset\mathcal{S}$$(x,y^{\star})\in\mathcal{B}$Sample on-policy response $\hat{y}\sim p_{S}(\cdot\mid x)$Compute the token-wise divergence along the student rollout:
@@
Calculate loss $\mathcal{L}_{\mathrm{OPSD}}(\theta)\leftarrow\frac{1}{|\mathcal{B}|}\sum_{(x,y^{\star})\in\mathcal{B}}\ell(x,y^{\star})$and update $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}_{\mathrm{OPSD}}(\theta)$
@@ 3.2 On-Policy Self-Distillation
Both policies share the same parameters $\theta$ but differ only in their conditioning context. To encourage the teacher to naturally evaluate the student’s generation, we add a prompt asking the teacher to generate a new solution after seeing the reference solution as shown in Figure 2. However, the teacher doesn’t generate tokens, it only does rationalization implicitly through prefilling.
@@
Following recent on-policy distillation methods (lu2025onpolicydistillation), we form a sampled-token reward signal (a reverse-KL signal on sampled actions) and optimize with policy gradient. For each position $n$ in a sampled sequence $\hat{y}$, define the advantage term
@@
OPSD as dense-reward policy gradient and comparison to STaR.
The objective in Equation 9 can be seen as policy gradient with dense, token-level rewards. In Appendix Appendix C, we formalize this and contrast with STaR (zelikman2022star), a closely related method that also uses the same model to generate reasoning traces, then performs rejection sampling followed by SFT on correct traces. This procedure can be viewed as policy gradient with a sequence-level binary reward that assigns identical credit to all tokens and vanishes when samples are incorrect. In contrast, OPSD provides feedback at every token position regardless of final-answer correctness.
[…]
Table 2: Performance comparison across mathematical reasoning benchmarks for Qwen3 models. We report average@16 using suggested sampling parameters from the Qwen3 blog with temperature of 1.0 and generation length of 38k, with detailed parameter in Table 6. For each method, we evaluate checkpoints every 100 training steps up to 1500 steps and report the best observed score within this training budget. This evaluation protocol is applied consistently across all baselines and we show accuracy over checkpoint steps in Figure 3. OPSD can match GRPO performance but not significantly outperform it, while OPSD can provide superior learning efficiency.
@@ 4 Experiments
We conduct comprehensive experiments to answer the following research questions:
@@
    How does OPSD compare to SFT and GRPO in reasoning performance and sample efficiency? (§4.2)
@@
    How does OPSD scale across model sizes? (§4.3.1)
@@
    What is the effect of generation length on performance and sample efficiency? (§4.3.2)
@@
    Does full-vocabulary divergence provide benefits over sampled-token policy gradient? (§4.3.3)
@@ 4.1 Experimental Setup
Baselines. We compare against two methods trained on the same dataset: (1) SFT, standard supervised fine-tuning on expert trajectories, which can be seen as off-policy distillation from a more powerful LLM that generated the reasoning traces; (2) GRPO (shao2024deepseekmath), group relative policy optimization with binary outcome rewards verified against ground-truth answers. The max generation length is set to 16k.
Implementation details. For GRPO, we sample 8 responses per problem. For OPSD, we sample 1 response per problem. We use Adam optimizer with a learning rate of 2e-5. For the divergence measure in Eq. 6, we use $\operatorname{JSD}_{\beta=0.5}$. We fix the teacher policy to be the initial policy, rather than the currently updating learning policy, as we find this helps stabilize training and implicitly acts as regularization to prevent excessive deviation from the initial policy. All experiments are conducted on 8×A100 GPUs with LoRA (hu2022lora). More experimental details are in Appendix B.
Implementation note on prompt formats. After the initial release of this preprint, an unintentional mismatch in the prompt format between the student’s sampling prompt and the teacher’s prompt was found: Specifically, the student prompt lacks the Qwen3 chat template, whereas the teacher prompts follow the Qwen3 thinking-style chat template. As a result, student reasonings were generated in a template-free style, while the teacher distributions were conditioned on prompts containing thinking tags. Although our method does not require the student and teacher prompt templates to match, we document this implementation detail here to ensure transparency and reproducibility. The experimental results reported in this preprint therefore correspond to this configuration. The prompt template/instructions can be viewed as a design choice to our method: both the teacher and student prompt formats can be treated as optimizable components of the framework. For instance, the teacher prompt can be refined via prompt optimization techniques such as GEPA (agrawal2025gepareflectivepromptevolution) to shape the supervising signal, while the student prompt can be varied to study which generation modes yield the most effective learning.
@@ 4.2 Main Results
Table 2 reports results on competition-level mathematical reasoning benchmarks. OPSD consistently outperforms SFT and improves over the base model across scales; it matches or exceeds GRPO at 4B/8B. However, we found OPSD to be under-performing at 1.7B scale, suggesting successful self-distillation requires suffcient model capability. Notably, OPSD accomplishes the gains using only a single rollout per problem, whereas GRPO requires 8 rollouts, demonstrating improved sample efficiency.
@@
OPSD is significantly more token-efficient than GRPO while matching the performance. Figure 3 compares the two methods under the same effective training batch size on Qwen3-4B. While GRPO relies on 8 rollouts with long generation budgets of 16k, OPSD achieves higher performance using substantially fewer generated tokens of 1024 and needs only 1 rollout per prompt. This efficiency stems from dense token-level supervision from the teacher distribution, reducing sampling cost and training time without sacrificing performance. We hypothesize that the early tokens are more important for distillation than the later tokens, as the earlier tokens can represent more important branching points.
@@ 4.3 Discussions
4.3.1 Effect of Model Scale
Our method relies on the teacher policy’s ability to rationalize reference solutions when conditioned on privileged information. Under a fixed dataset, this capability depends on sufficient model capacity and is expected to scale with model size. We therefore hypothesize that OPSD becomes increasingly effective as models grow more capable of leveraging privileged context. To evaluate this, we apply OPSD to the Qwen3 family at three scales: 1.7B, 4B, and 8B parameters. As shown in Table 2, OPSD yields larger improvements at the 4B and 8B scales than at the 1.7B scale. While OPSD improves performance of the 1.7B model on AIME24, AIME25, and AMO-Bench, it decreases performance on HMMT25. We hypothesize that the learning signal may be noisy when the student model is unable to effectively utilize the provided privileged information, and that the underlying tasks may be too challenging for a model of this capacity, consistent with our hypothesis that self-distillation relies on sufficient model ability.
4.3.2 Effect of Generation Length
@@ 5 Related Work
LLM Self-Training.
Our work connects to a line of research showing that LLMs can improve by generating and exploiting their own supervision signals (allentowards; xu2024survey; chen2024self; wang2023self; sun2023principle; yuan2024self; yang2024self). Closest in spirit is context distillation (snell2022learning), which uses the same underlying model as both teacher and student by providing the teacher with privileged context and then SFT the student on the teacher’s generated outputs without context. This can be viewed as off-policy, where the learning signal is a discrete token sequence. In the reasoning domain, ReST (gulcehre2023reinforced) and STaR (zelikman2022star) similarly rely on iterative self-training loops—generate rationales conditioned on hints or answers, filter by rewards or ground-truth answers, and fine-tune on successful trajectories—again yielding hard distillation; mitra2025semantic extends this to soft distillation. In-context editing (qicontext) does on-policy sample from student and shows that context-induced knowledge can be internalized via soft distillation by minimizing divergences and demonstrates this in knowledge editing settings. OPSD differs from these approaches in that we perform on-policy, soft distillation on the student’s own rollouts for reasoning tasks: the teacher’s supervision is per-token distribution matching rather than generating a rationale for SFT. OPSD frames reasoning improvement as learning a conditional distribution induced jointly by the dataset’s ground-truth solutions and the model’s own reasoning ability. Concurrently, SDPO (hubotter2026reinforcement) explored similar algorithm with environment feedbacks as privilledged information and SDFT (shenfeld2026selfdistillationenablescontinuallearning) explored on-policy self-distillation on continual learning tasks.
@@
Improving LLM Reasoning through SFT and RL. SFT and RL are two primary methods for improving LLM reasoning ability. SFT on high-quality reasoning traces has demonstrated strong performance (yu2023metamath; numina_math_datasets; pasteropenwebmath; openthoughts; ye2025limoreasoning; muennighoff2025s1; zhou2023lima). However, prior work shows that SFT can rely on memorization rather than robust generalization (chu2025sft). In contrast, RL optimizes directly for outcome-based objectives can exhibit better generalization (huan2025does). More recent algorithms such as GRPO (guo2025deepseek; shao2024deepseekmath) enable scalable RL by estimating advantages from group-level rewards without requiring an explicit critic as in PPO (schulman2017proximal). Building on this line of work, a growing body of research highlights the effectiveness of RLVR for reasoning tasks (yu2025dapo; liu2025understanding; yue2025vapo; Polaris2025; zheng2025group).
@@ 6 Conclusion
7 Impact Statement
This paper presents work whose goal is to advance the field of machine learning. Our method improves the efficiency of training language models for reasoning tasks, reducing computational costs compared to existing reinforcement learning approaches. We do not foresee specific negative societal consequences.
Appendix A Limitations and Future Directions
@@
Appendix B Experimental Details
We provide the training and evaluation configurations for our SFT, GRPO and OPSD experiments in Tables 5, 4 and 6.
Evaluation protocols in Table 2: For each method (GRPO, OPSD and SFT), we report the highest achieved score obtained by evaluating every 100 steps of training, up to 1500 steps, for each task in our results. Since we apply the same protocol consistently to all baselines, this remains a fair comparison across different algorithms given the inherent variance at any fixed evaluation point. We also report averaged accuracy over checkpoints for all three baselines in Figure 3. In Figure 3, OPSD runs with generation length of 1024 and learning rate of $2\times 10^{-4}$. For OPSD, we find $2\times 10^{-4}$ yield better results than with $2\times 10^{-5}$ for lower generation length like 1024.
Table 4: Training Configuration for GRPO and OPSD
Parameter GRPO OPSD
Learning Rate $2\times 10^{-5}$ $2\times 10^{-5}$ / $2\times 10^{-4}$
Batch Size (per device) 1 1
Gradient Accumulation Steps 4 4
Effective Batch Size 32 32
LoRA Rank ($r$) 64 64
LoRA Alpha ($\alpha$) 128 128
LoRA Target Modules q_proj, k_proj, v_proj, o_proj,
gate_proj, up_proj, down_proj
Max Completion Length 16,000 2048/1024
Number of Generations per Prompt 8 1
Sampling Temperature 1.2 1.1
KL Coefficient ($\beta$) 0.0
Table 5: Training Configuration for SFT.
Parameter SFT
Learning Rate $2\times 10^{-5}$
Batch Size (per device) 2
Gradient Accumulation Steps 4
Effective Batch Size 64
LoRA Rank ($r$) 64
LoRA Alpha ($\alpha$) 128
LoRA Target Modules q_proj, k_proj, v_proj, o_proj,
gate_proj, up_proj, down_proj
Max Sequence Length 16000
Number of Training Epochs 4
Training Dataset Size 30k
Table 6: Evaluation Parameters.
Parameter Value
Max New Tokens 38912
Thinking Mode Enabled
Top-p 0.95
Top-k -1
Min-p 0.0
Presence Penalty 0.0
Samples per Prompt 16
Temperature 1.0
All experiments were conducted using 8 A100 or 8 H100 GPUs with gradient checkpointing and Flash Attention 2 for memory efficiency. We use the AdamW (loshchilov2017decoupled) optimizer and bfloat16 precision for all training runs. For OPSD, unless otherwise stated, we used full-vocabulary logit distillation.
Appendix C Policy-Gradient Interpretation of OPSD and Comparison to STaR
Our OPSD objective in Equation 9 can be interpreted as a policy-gradient update with a dense, token-level reward signal derived from privileged information. In this section, we show: (1) OPSD can be seen as a dense-reward policy gradient, and (2) we contrast OPSD with STaR, demonstrating that STaR’s learning signal is sequence-level while OPSD is token-level.
C.1 STaR as Sequence-Level Policy-Gradient
STaR (zelikman2022star) can be viewed as an approximation to an RL-style policy gradient objective. The language model $p_{\theta}$ induces a joint distribution over rationale $r$ and answer $y$:
$p_{\theta}(r,y\mid x)=p_{\theta}(r\mid x)\,p_{\theta}(y\mid x,r),$
where the model first samples a latent rationale $r$ before predicting the final answer $y$. Given an indicator reward $R(y)=\mathbf{1}(y=y^{\star})$, the expected return across the dataset $\mathcal{S}=\{(x_{i},y_{i}^{\star})\}_{i=1}^{N}$ is
$J_{\text{STaR}}(\theta)=\sum_{i=1}^{N}\mathbb{E}_{(r,y)\sim p_{\theta}(\cdot\mid x_{i})}\big[\mathbf{1}(y=y_{i}^{\star})\big].$ (10)
Applying the log-derivative trick yields a policy gradient:
$\nabla_{\theta}J_{\text{STaR}}(\theta)=\sum_{i=1}^{N}\mathbb{E}_{(r,y)\sim p_{\theta}(\cdot\mid x_{i})}\Big[\mathbf{1}(y=y_{i}^{\star})\,\nabla_{\theta}\log p_{\theta}(r,y\mid x_{i})\Big].$ (11)
Note that the indicator function discards the gradient for all sampled rationales that do not lead to the correct answer $y_{i}^{\star}$: this corresponds to the filtering step in STaR.
One limitation is that STaR’s reward is sequence-level: the binary indicator $\mathbf{1}(y=y^{\star})$ provides the same signal to all tokens in a trajectory, offering no intermediate credit assignment. When all sampled trajectories are all incorrect, the learning signal vanishes.
C.2 OPSD as Dense-Reward Policy Gradient
The sampled-token objective in Equation 9 can also be viewed as a policy-gradient method, but with a token-level reward. Fix a training pair $(x,y^{\star})$ and let the student generate a trajectory $\hat{y}\sim p_{S}(\cdot\mid x)$. At each position $n$, define the per-token reward:
$r_{n}(x,\hat{y})\triangleq\log p_{T}(\hat{y}_{n}\mid x,y^{\star},\hat{y}_{<n})-\log p_{S}(\hat{y}_{n}\mid x,\hat{y}_{<n}).$
This reward measures how much the privileged teacher prefers the sampled token $\hat{y}_{n}$ relative to the student. As stated in the main text, we treat $r_{n}$ (equivalently, the advantage $A_{n}$) as a constant with respect to $\theta$ when computing gradients—that is, we stop gradients through both $p_{T}$ and $p_{S}$ in the reward computation. Under this treatment, the gradient of Equation 9 takes the standard policy-gradient form:
$\nabla_{\theta}\mathcal{L}(\theta)=-\mathbb{E}_{(x,y^{\star})\sim\mathcal{S}}\left[\mathbb{E}_{\hat{y}\sim p_{S}(\cdot\mid x)}\left[\frac{1}{|\hat{y}|}\sum_{n=1}^{|\hat{y}|}r_{n}(x,\hat{y})\,\nabla_{\theta}\log p_{S}(\hat{y}_{n}\mid x,\hat{y}_{<n})\right]\right],$
which corresponds to maximizing the expected per-token reward along on-policy student rollouts:
[…]
This reward is dense: it provides a learning signal at every token position, regardless of whether the final answer is correct.
Comparison.
Both STaR and OPSD can be understood as policy-gradient methods, but their reward structures differ fundamentally. STaR uses a sequence-level indicator $\mathbf{1}(y=y^{\star})$ that assigns the same signal to all tokens; when all sampled trajectories are incorrect, the learning signal vanishes entirely. In contrast, OPSD provides a token-level reward $r_{n}$ at every position, enabling fine-grained credit assignment even when the final answer is wrong.